OpenStudy (gabylovesyou):

How can one halfx − 5 = one thirdx + 6 be set up as a system of equations ?

OpenStudy (gabylovesyou):

\[\frac{ 1 }{ 2 }x - 5 =\frac{ 1 }{ 3 }x + 6\]

OpenStudy (xmissalycatx):

Do you know the formula for a system of equations?

OpenStudy (gabylovesyou):

I know how to solve as in substitution, elimination.. stuff like that.. if its something else, then no idk.

OpenStudy (mathmale):

A better question might be: Are you familiar with systems of linear equations? If so, type out an example of a system of linear equations (with 2 equatins, 2 variables).

OpenStudy (xmissalycatx):

You want that equation above to be in x + y = (a number)

OpenStudy (gabylovesyou):

y = 3x-2 y = -x - 6 example.

OpenStudy (xmissalycatx):

If it is a linear equation, you want it to be y = mx + b form..

OpenStudy (astrophysics):

Yeah so basically you have two equation that are equal to each other, what I mean is they have it set up as y = y: \[y=1/2x-5\] and \[y=1/2x+6\] I mean I was going to give you a different example but I think it's a bit obvious in either case.

OpenStudy (xmissalycatx):

Pretty much ^

OpenStudy (gabylovesyou):

my choices are.. A. 2y + x = −10 3y + x = 18 B. 2y + 2x = −10 3y + 3x = 18 C. 2y − x = −10 3y − x = 18 D. 2y − 2x = −10 3y − 3x = 18

OpenStudy (astrophysics):

It's the same thing, requires a little algebra, mess around with it and I think you will see :-)

OpenStudy (mathmale):

\[\frac{ 1 }{ 2 }x - 5 =\frac{ 1 }{ 3 }x + 6\] ... could be re-written in the equivalent form \[\frac{ 1 }{ 2 }x - 5 =y=y~=\frac{ 1 }{ 3 }x + 6\]

OpenStudy (mathmale):

from which we get the system of linear equations\[y=(1/2)x-5,~y=(1/3)x+6\]

OpenStudy (mathmale):

or \[y=(1/2)x-5\]

OpenStudy (mathmale):

\[y=(1/3)x+6\]

OpenStudy (gabylovesyou):

the answer is C... correct ?

OpenStudy (astrophysics):

Looks good!